If you have 0.20 mol of silver permanganate (AgMNO4), which has a solubility of about 16.7...

  1. If you have 0.20 mol of silver permanganate (AgMNO4), which has a solubility of about 16.7 g/L, what is the minimum amount of water you’ll need to get it all into solution? At what temperature? Report your answer to the nearest whole mL.
Answer #1

Given solubility 16.7 g/L is at 25oC hence the calculated amount of water is also at 25oC.

The molar mass of AgMnO4 = 226.8 g/mol

Mass of 0.20 mole of AgMnO4 = 0.20*226.8 = 45.36 g

Thus the minimum amount required to make a solution = 45.36g / 16.7 g/L = 2.716 L = 2716 mL

Thus the minimum amount of water needed is = 2716 mL.